Circular No. 7931

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 L9 (NEAT) An object originally reported as asteroidal on NEAT CCD images

taken at Palomar was posted on the NEO Confirmation Page due to its unusual orbit. S. Pravdo later suggested the object was cometary, in particular on July 2 noting a central condensation of diameter 3′′ surrounded by a spherical nebulosity of diameter ∼ 8′′. Discovery observation:

2002 UT  α_{2000}  δ_{2000}  m_{2}
June 13.40771  20^{h}02^{m}53.^{s}28  −27^{o}46′20′′.0  17.6

Available astrometry (with June 6 prediscovery observations), the following orbital elements, and an ephemeris appear on MPEC 2002-N15.

T = 2004 Apr. 10.967 TT  ω = 232.^{o}066
Ω = 110.433 }2000.0
q = 6.98834 AU  i = 68.251

COMETS C/1997 O2, C/1997 S4, C/1997 U8, C/1997 U9 (SOHO) Further to IAUC 7930, D. Hammer reports his measurements for ad-

ditional comets found on SOHO website C2 images by R. Kracht. Reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. Comet C/1997 S4 does not appear to belong to any of the established groups; the others belong to the Meyer group (cf. IAUC 7832).

Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 O2 July 25.313  8^{h}15.^{m}1  +20^{o}39′  2002-M23
C/1997 S4 Sept.29.188  12 16.7  − 1 08  2002-M47
C/1997 U8 Oct. 19.769  13 31.5  − 9 11  2002-M43
C/1997 U9  23.024  13 45.9  − 9 18  2002-M47

COMETS C/2001 J5 AND C/2001 K10 (SOHO) Hammer also reports measurements for additional Kreutz sungrazing

comets found on SOHO website C2 images by X. Leprette. Reduced astrometry and orbital elements by Marsden appear on MPEC 2002-M43.

Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 J5  May 13.112  3^{h}23.^{m}7  +17^{o}00′
C/2001 K10  18.936  3 45.9  +18 17
2002 July 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green